OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12 (Points : 1) Question 5. 5. What is the slope of the

OpenStudy (anonymous):

Well slope intercept is always in terms of y. since y = mx + b... Therefore, isolate y on the left side. y = (8/6)x + 2

OpenStudy (anonymous):

Slope is \[\frac{ dy }{ dx }\] or the rise over run... delta y/ delta x Since this graph is linear, the slope is infront of the x term. Which is 8/6